A mortgage of 665,000 is amortized over 25 years by makind end of month payments of 3560.



What is the annual rate of interest compounded semi annualy?



What is value of final payment?



What is the total interest cost for the debt over 25 years?

1
Expert's answer
2021-12-20T19:06:41-0500

What is the annual rate of interest compounded semi annually?

Yearly payments;

=monthly payments "\times" 12

"=3560\times12=42720"

annual rate of interest

"A=P[(1+r)^n-1]"

"42720=665000[(1+r)^n-1]"

"(1+r)^n=\frac{42720}{665000}+1"

"(1+r)^n=1.0642"

Since n=1

"r=1.064-1=0.064=6.4\%"

Therefore annual rate of interest compounded semi annually is;

"=\frac{r}{2}=\frac{6.4\%}{2}=3.2\%"

What is value of final payment?

=yearly payments"\times" number of years

"=42720\times25"

"=1,068,000"


What is the total interest cost for the debt over 25 years?

Interest="1,068,000-665000=403,000" '

Interest rate="\frac{403,000}{665000}=0.606=60.6\%"